OpenStudy (anonymous):

(square root of t-2) /(3t-9)

hero (hero):

\[\frac{\sqrt{t-2}}{3t-9}\]

OpenStudy (anonymous):

yes need the domain of it.

hero (hero):

@Zelda, the domain is limited to more than just that

OpenStudy (anonymous):

Oops, sorry.

OpenStudy (kyosuke):

No es cualquier cosa,como parte de la función es racional "t es diferente de 3" ademas t-2>= 0 ; eso es el dominio. suerte!.

hero (hero):

Hint: Solve t - 2 > 0 and 3t - 9 ≠ 0

OpenStudy (anonymous):

t=2 and t does not equal to 3

hero (hero):

sorry, t ≥ 2 and t ≠ 3 or [2, 3) (3, ∞)

OpenStudy (kyosuke):

@Hero "t" si es Igual a 2, el Dom f = [2, ∞), t ≠ 3

hero (hero):

I don't see any difference in what you wrote from what I wrote

OpenStudy (kyosuke):

pusiste t ≠ 3 o 2

hero (hero):

Mind saying that in English?

OpenStudy (kyosuke):

no entiendo

hero (hero):

\(\sqrt{0} = 0\)

hero (hero):

\(\sqrt{2 - 2} = 0\)

OpenStudy (kyosuke):

que con eso